out of 2000 families with 4 children, how many would you expect to have 1 or 2 girls, given that the probability of a male birth is 0.5____
1
Expert's answer
2017-06-03T05:02:09-0400
The probability of having 1 girl: P(X=1)=C_4^1 0.5^1 0.5^3=0.25.
The probability of having 2 girls: P(X=2)=C_4^2 0.5^2 0.5^2=0.375.
The probability of having 1 or girls: P=0.25+0.375=0.625.

Out of 2000 families with 4 children N=2000*P=2000*0.625=1250 are expected to have 1 or 2 girls.